OpenStudy (anonymous):

From a group of 6 women and 8 men a committee consisting of 3 men and 4 women is to be formed. How many different committees are possible if (a) 2 of the men refuse to serve together? (b) 2 of the women refuse to serve together? (c) 1 man and 1 woman refuse to serve together?

OpenStudy (anonymous):

I already have the answer for part c, it's 630. Help with a and b please?

OpenStudy (anonymous):

For (a) Let name the men as \(A_1,A_2,\cdots A_8 \) and \( A_1 \) and \( A_2\) don't want to serve together. Now you will have two cases. (1) Taking either of \( A_1 \) or \( A_2\) from those two men and rest two from the remaining 6 men. \[ \binom 2 1 \times \binom 6 2 \times \binom 8 4 \] (2) Taking the three men from 6 men. \[ \binom 6 3 \times \binom 8 4 \] Your answer is sum of these two cases. For (b) the idea is practically same.

OpenStudy (anonymous):

wazzup

OpenStudy (anonymous):

Thank you so much :) :)

OpenStudy (anonymous):

Glad to help :)